OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, nd the limit n!/2^n

OpenStudy (anonymous):

You can use the ratio test. \[\lim_{n->\infty}\left|\frac{a_{n+1}}{a_n} \right|=\lim_{n->\infty}\frac{(n+1)!/2^{n+1}}{n!/2^{n}}=\lim_{n->\infty}\frac{(n+1)n!}{2 \times 2^n}\frac{2^n}{n!}\]\[=\lim_{n->\infty}\frac{n}{2} \rightarrow \infty\] The sequence does not converge since the ratio in the limit is not less than 1.

OpenStudy (anonymous):

nice!

OpenStudy (anonymous):

Thanks ;)

OpenStudy (anonymous):

\[\frac{n+1}{2}\]

OpenStudy (anonymous):

A question.. So does it mean that you cant test some values as N... and see whats happening?

OpenStudy (anonymous):

Can you rephrase that? I'm not sure what you mean.

OpenStudy (anonymous):

n!/2^n... Put in some values, as a check.. ..n= 1, 2, 3

OpenStudy (anonymous):

And see if it diverges or converges

OpenStudy (anonymous):

I hope you didnt misunderstand me, it was just a question, if i would in some values, in the funktion , n!/2^n .. could i see right away if it diverges or converges..

OpenStudy (anonymous):

Ah I see. No, because that's not actually testing for ALL n. The tests are derived from the definition of convergence.

OpenStudy (anonymous):

You could get an idea, but it wouldn't prove it. That's all.

OpenStudy (anonymous):

Okey, so u have to, show it for all n. like u did

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

Well, thx.

OpenStudy (anonymous):

No worries.

OpenStudy (anonymous):

Hehe, I thought this was XiaoHong's question.

OpenStudy (anonymous):

So did i ;)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome.